OpenStudy (anonymous):

X - 2 = sqrt(3) If I want to square root both sides, everything has to be squared right?

OpenStudy (anonymous):

Correction: If I want to SQUARE both sides

OpenStudy (shubhamsrg):

yes..both sides will be squared..

OpenStudy (anonymous):

@shubhamsrg So (x-2)^2 = (sqrt(3))^2 right?

OpenStudy (shubhamsrg):

yep..the very same..

OpenStudy (anonymous):

Squaring both sides will leave you with a quadratic equation. It'll become exactly as you've posted above. On simplifying, you get, x^2 + 4 - 4x = 3 x^2 - 4x -1 = 0

OpenStudy (anonymous):

But everyone has to be squared right? So can I image everything on the sides of the equations being encompassed by parentheses then raised to the power of 2?

OpenStudy (anonymous):

Example: 2x + 3y + 2 = 3 + 7 Squaring both sides (2x + 3y + 2)^2 = (3 + 7)^2

OpenStudy (shubhamsrg):

yes..you can imagine that ;)

OpenStudy (shubhamsrg):

yep..

OpenStudy (anonymous):

Thanks

OpenStudy (shubhamsrg):

glad to help ^_^

OpenStudy (shubhamsrg):

but mind you if you square some eqn,,for eg. in the eqn you had posted.. it was converted into a quadratic eqn..if you solve it..you'll get 2 values of x.. both will not satisfy the original eqn! so be careful..

OpenStudy (shubhamsrg):

this is ought to happen,,as its like this when x-2 =0 and you multiply both sides by (x-3) and then solve,,you get x=2 and 3,,but you know only 2 satisfies original one.. so you see..why it happens..